# Lowering PyTorch’s Memory Consumption for Selective Differentiation

### v1

@@ 3.1 ResNet-101
For ResNet-101, we take a detailed look at our memory-saving layer’s effects. It is a decently powerful, modern model and frequently used as a backbone for other architectures such as CLIP [25] and LAVA [11]. It contains dense, convolution, batch normalization, ReLU, and max/average pooling layers. Table 2 summarizes our findings.
PyTorch is unaware of selective differentiation:
In the upper part of Table 2, we see that the default PyTorch implementation uses the same amount of memory for any scenario. In fact, marking only the input to a neural net as differentiable consumes as much memory as marking all parameters, although the former does not require storing the inputs to linear layers. This confirms the findings on the toy model from Section 2.
@@
After swapping out ReLU with our mask-based custom implementation, we observe substantial memory savings. E.g., memory consumption for ‘Input’ dropped by a factor of almost 5x on GPU. Only when all parameters are differentiable, we see a slight increase in memory after swapping out ReLUs. This is to be expected as the mask stored by our custom implementation requires additional storage.
@@
3.2 More results
To further solidify our findings, we now evaluate our layers on other popular and commonly used CNNs, including ResNet-18 [13], VGG-16 [29], ResNeXT101-64x4d [35], EfficientNetv2-L [30, 31], MobileNetv3-L [28, 14], FCN with a ResNet-101 backbone [20], DeepLabv3 with a ResNet-101 backbone [4], Faster-RCNN with a ResNet-50 backbone [27], and SSDLite with a MobileNetv3-L backbone [28]. Table 3 summarizes the comparison.
On this large repertoire of networks, we observe the same effects as on ResNet-101 from the previous section: Due to our customized ReLU implementation, memory is slightly higher in the absence of selective differentiation. Run time is unaffected by swapping in our layers, while the selective differentiation scenarios consistently show lower memory consumption with reduction factors up to 6x, underlining the usefulness of our approach in this context.
@@ 4 Conclusion
We have shown how to improve the memory consumption of PyTorch’s automatic differentiation in the context of selective differentiation where gradients are only requested for a subset of variables—a common situation in modern fine-tuning tasks. To overcome this, we provide a drop-in implementation which takes into account the differentiability of all tensors when storing the computation graph. Empirically, we demonstrated the effectiveness of our approach to reduce memory in multiple selective differentiation cases without affecting run time.
Our method is easy to use, requiring only a single call to a converter function that replaces all supported layers with our equivalents. Next, we plan to apply it to parameter-efficient LLM fine-tuning or layer norm tuning, and other modern architectures like vision transformers. We expect this to require additional efforts due to the interaction with other transformer layers like softmax and GeLU.

### v2

@@ 3.1 ResNet-101
For ResNet-101, we take a detailed look at our memory-saving layer’s effects. It is a decently powerful, modern model and frequently used as a backbone for other architectures such as CLIP Radford et al. (2021) and LAVA Gurram et al. (2022). It contains dense, convolution, batch normalization, ReLU, and max/average pooling layers. Table 2 summarizes our findings (model in training mode).
PyTorch is often unaware of selective differentiation:
In the upper part of Table 2, we see that the default PyTorch implementation uses the same amount of memory for any scenario. In fact, marking only the input to a neural net as differentiable consumes as much memory as marking all parameters, although the former does not require storing the inputs to fully-connected and convolution layers. This confirms the findings on the toy model from Section 2.
@@
After swapping out ReLU with our mask-based custom implementation, we observe substantial memory savings. E.g., memory consumption for ‘Input’ dropped to roughly two thirds. Only when all parameters are differentiable, we see a slight increase in memory after swapping out ReLUs. This is to be expected as the mask stored by our custom implementation requires additional storage.
@@
3.2 Results on CNNs
To further solidify our findings, we now evaluate our layers on other popular and commonly used CNNs, including ResNet-18 He et al. (2016), VGG-16 Simonyan & Zisserman (2015), ResNeXt101-64x4d Xie et al. (2017), EfficientNetv2-L Tan & Le (2019; 2021), MobileNetv3-L Sandler et al. (2018); Howard et al. (2019), FCN with a ResNet-101 backbone Long et al. (2015), DeepLabv3 with a ResNet-101 backbone Chen et al. (2017), Faster-RCNN with a ResNet-50 backbone Ren et al. (2015), and SSDLite with a MobileNetv3-L backbone Sandler et al. (2018). Table 3 summarizes the comparison.
On this large repertoire of networks, we observe the same effects as on ResNet-101 from the previous section: Due to our customized ReLU implementation, memory is slightly higher in the absence of selective differentiation. Run time is unaffected by swapping in our layers, while the selective differentiation scenarios consistently show lower memory consumption (as low as two thirds), underlining the usefulness of our approach in this context.
So far, we used all batch normalization layers in training mode. In Appendix A, we experiment with BatchNorm2d in evaluation mode. This allows discarding normalization layer inputs whenever the parameters are marked non-differentiable and enables further memory savings, with reductions up to 6x and no runtime overhead (Table 5).
3.3 Results on Transformers
Memory [GiB] Time [s]
Case All Input Norm Surgical All Input Norm Surgical
SiLU Transformers
LLaMa3-8B Touvron et al. (2023) ($H$ = 4096) 31.01 (1.00) 27.27 (0.88) 28.26 (0.91) 28.18 (0.91) 1.39 (1.00) 1.13 (0.81) 1.13 (0.82) 1.17 (0.84)
+ MemSave 29.01 (0.94) 26.26 (0.85) 26.26 (0.85) 26.94 (0.87) 1.61 (1.16) 1.04 (0.75) 1.05 (0.75) 1.12 (0.81)
Mistral-7B Jiang et al. (2023) ($H$ = 4096) 41.67 (1.00) 34.20 (0.82) 36.17 (0.87) 36.01 (0.86) 2.09 (1.00) 1.55 (0.74) 1.56 (0.75) 1.67 (0.80)
+ MemSave 37.67 (0.90) 32.17 (0.77) 32.17 (0.77) 33.54 (0.80) 2.55 (1.22) 1.44 (0.69) 1.46 (0.70) 1.71 (0.82)
Phi3-4B Gunasekar et al. (2023) ($H$ = 3072) 31.74 (1.00) 26.01 (0.82) 27.49 (0.87) 27.40 (0.86) 1.59 (1.00) 1.23 (0.78) 1.24 (0.78) 1.31 (0.83)
+ MemSave 28.74 (0.91) 24.49 (0.77) 24.49 (0.77) 25.55 (0.81) 1.69 (1.06) 1.09 (0.69) 1.12 (0.71) 1.23 (0.78)
ReLU Transformers
Transformer Vaswani et al. (2017) ($H$ = 2048) 26.91 (1.00) 21.54 (0.80) 21.54 (0.80) 23.04 (0.86) 2.47 (1.00) 2.21 (0.90) 2.19 (0.89) 2.24 (0.91)
+ MemSave 25.60 (0.95) 20.23 (0.75) 20.23 (0.75) 21.73 (0.81) 2.57 (1.04) 2.26 (0.92) 2.25 (0.91) 2.30 (0.93)
T5 Raffel et al. (2020) ($H$ = 768) 33.40 (1.00) 25.94 (0.78) 28.85 (0.86) 27.77 (0.83) 1.70 (1.00) 1.37 (0.81) 1.39 (0.82) 1.47 (0.86)
+ MemSave 31.84 (0.95) 22.80 (0.68) 22.80 (0.68) 25.28 (0.76) 1.95 (1.14) 1.52 (0.89) 1.54 (0.90) 1.59 (0.94)
Table 4: GPU peak memory and run time comparison between PyTorch and our memory-saving layers for LLMs.
Selective differentiation is of high practical relevance for large language models (LLMs), especially with the increasing popularity of vision language models (VLMs) and the release of models such as LLaVa, PaLI, PaLI-Gemma and GPT-4o Liu et al. (2023); Chen et al. (2023b; a). Another area where this is quite important is using a (light) modality-specific encoder to encode (and project) the input of different modalities into the embedding space of the LLM, such as in scientific ML Shen et al. (2024).
Here, we investigate the impact of our memory-saving layers on attention-based models. Popular libraries like HuggingFace that provide access to such models often implement attention through linear layers, which we did not find to suffer from the behavior of convolutions. However, an additional challenge in these architectures is dropout, which may interact with inputs to linear layers similarly to other activations (Section 2.2.2). Thus, we find our tricks for ReLU and dropout helpful to avoid storing inputs to linear layers if their weights are marked non-differentiable.
We consider the same cases as in Section 3 and all nets consume embeddings of size (64, 256, $H$), with hidden size $H$ of a network (given in Table 4). We test on the vanilla Transformer Vaswani et al. (2017), T5 Raffel et al. (2020), LLaMa3 Touvron et al. (2023), Mistral Jiang et al. (2023) and Phi3 Abdin et al. (2024); Gunasekar et al. (2023). The results are shown in Table 4. With LLaMa (8B parameters), Mistral (7B parameters) and Phi3 (4B parameters) being large models, they are loaded in bfloat16 data type and the batch size is reduced to 8 for LLaMa and 16 for the others. We observe that our approach also enables memory savings on attention-based models and consistently achieves a smaller memory footprint.
@@ 4 Conclusion
We have shown how to improve the memory consumption of PyTorch’s automatic differentiation in the context of selective differentiation where gradients are only requested for a subset of variables—a common situation in modern fine-tuning tasks. Our approach is based on the insight that PyTorch sometimes ignores the differentiability of tensors; specifically in (transpose) convolutions and batch normalization layers in evaluation mode. To overcome this, we provide a drop-in implementation which takes into account the differentiability of all tensors when storing the computation graph. Empirically, we demonstrated the effectiveness of our approach to reduce memory in multiple selective differentiation cases without affecting run time, on both convolution- and attention-based architectures. Our method is easy to use, requiring only a single call to a converter function that replaces all supported layers with our equivalents. Next, we plan to study its impact on real-world applications, such as parameter-efficient LLM fine-tuning with low-rank adapters Hu et al. (2022).
@@ Acknowledgements
Appendix A Results on CNNs when BatchNorm layers are set to eval mode
Memory [GiB] Time [s]
case All Input Norm SurgicalFirst All Input Norm SurgicalFirst
DeepLabv3 (RN101) Chen et al. (2017) 22.82 (1.00) 22.82 (1.00) 22.82 (1.00) 22.82 (1.00) 0.88 (1.00) 0.66 (0.75) 0.70 (0.79) 0.70 (0.79)
+ MemSave 24.90 (1.09) 4.28 (0.19) 15.17 (0.66) 8.40 (0.37) 0.90 (1.02) 0.69 (0.79) 0.72 (0.82) 0.73 (0.83)
EfficientNetv2-L Tan & Le (2019; 2021) 26.81 (1.00) 26.81 (1.00) 26.81 (1.00) 26.81 (1.00) 0.74 (1.00) 0.59 (0.79) 0.59 (0.80) 0.64 (0.86)
+ MemSave 26.81 (1.00) 10.45 (0.39) 18.64 (0.70) 17.31 (0.65) 0.75 (1.01) 0.58 (0.78) 0.59 (0.80) 0.66 (0.89)
FCN (RN101) Long et al. (2015) 22.23 (1.00) 22.23 (1.00) 22.23 (1.00) 22.23 (1.00) 0.79 (1.00) 0.59 (0.75) 0.62 (0.79) 0.63 (0.80)
+ MemSave 24.39 (1.10) 4.26 (0.19) 15.15 (0.68) 7.99 (0.36) 0.81 (1.03) 0.63 (0.79) 0.65 (0.82) 0.67 (0.85)
Faster-RCNN (RN101) Ren et al. (2015) 6.84 (1.00) 6.84 (1.00) 6.84 (1.00) 6.84 (1.00) 0.75 (1.00) 0.64 (0.84) 0.64 (0.85) 0.67 (0.88)
+ MemSave 7.31 (1.07) 1.98 (0.29) 4.79 (0.70) 4.19 (0.61) 0.74 (0.99) 0.63 (0.83) 0.63 (0.83) 0.66 (0.87)
MobileNetv3-L Howard et al. (2019) 2.82 (1.00) 2.82 (1.00) 2.82 (1.00) 2.82 (1.00) 0.40 (1.00) 0.31 (0.78) 0.31 (0.78) 0.34 (0.86)
+ MemSave 2.96 (1.05) 0.87 (0.31) 1.91 (0.68) 2.10 (0.75) 0.39 (0.98) 0.31 (0.79) 0.31 (0.79) 0.35 (0.87)
ResNeXt101-64x4d Xie et al. (2017) 15.15 (1.00) 15.15 (1.00) 15.15 (1.00) 15.15 (1.00) 0.62 (1.00) 0.48 (0.77) 0.50 (0.80) 0.53 (0.85)
+ MemSave 16.75 (1.11) 2.46 (0.16) 9.87 (0.65) 9.77 (0.64) 0.61 (0.99) 0.47 (0.77) 0.49 (0.79) 0.52 (0.84)
SSDLite (MobileNetv3-L) Sandler et al. (2018) 0.54 (1.00) 0.53 (0.97) 0.53 (0.97) 0.53 (0.97) 0.62 (1.00) 0.56 (0.90) 0.52 (0.84) 0.57 (0.91)
+ MemSave 0.57 (1.04) 0.26 (0.48) 0.41 (0.75) 0.44 (0.82) 0.63 (1.02) 0.58 (0.93) 0.53 (0.86) 0.57 (0.92)
Table 5: GPU peak memory and run time comparison between PyTorch and our memory-saving layers for CNNs with BatchNorm layers in eval mode. VGG-16 has been excluded here as it does not contain any BatchNorm layers.
Appendix B Probing PyTorch Layers in the Presence of Selective Differentiation
Our experiments with a deep CNN of size-preserving convolutions from Figures 1 and 2.1 revealed that PyTorch’s 2d convolution layer stores its input tensor whenever it is differentiable, irrespective of the weight’s differentiability. Here, we perform analogous experiments, but with other layers we suspect to exhibit similar behaviour. Specifically, we consider layers whose forward pass is linear w.r.t. both the layer input and the layer weights. This includes fully-connected layers ( torch.nn.Linear), convolution layers ( torch.nn.ConvNd), transpose convolution layers ( torch.nn.ConvTransposeNd), and batch normalization layers in evaluation mode (e.g.  torch.nn.BatchNorm2d).
[…]
  • $(512,1024,256)$ for torch.nn.Linear
  • $(4096,8,4096)$ for torch.nn.Conv1d and torch.nn.ConvTranspose1d
  • $(512,8,256,256)$ for torch.nn.BatchNorm2d, torch.nn.Conv2d, and torch.nn.ConvTranspose2d
  • $(64,8,64,64,64)$ for torch.nn.Conv3d and torch.nn.ConvTranspose3d
Compared to the input and intermediate activations, the memory footprint of a layer’s weight is negligible:
  • $(256,256)$ for torch.nn.Linear
  • $(8,8,3)$ for torch.nn.Conv1d and torch.nn.ConvTranspose1d
  • $(8)$ for torch.nn.BatchNorm2d
  • $(8,8,3,3)$ for torch.nn.Conv2d and torch.nn.ConvTranspose2d
  • $(8,8,3,3,3)$ for torch.nn.Conv3d and torch.nn.ConvTranspose3d
Our results are summarized in Figure 4 without using torch.compile, and in Figure 6 with torch.compile. In both scenarios, we can see that PyTorch’s convolutions, transpose convolutions, and batch normalization (in evaluation mode) are not agnostic to the differentiability of their weights. PyTorch’s fully-connected layer, however, is already agnostic to the differentiability of its weights.
Training mode
Evaluation mode
(a) torch.nn.Linear
(b) torch.nn.Conv1d
(c) torch.nn.Conv2d
(d) torch.nn.Conv3d
(a) torch.nn.ConvTranspose1d
(b) torch.nn.ConvTranspose2d
(c) torch.nn.ConvTranspose3d
(d) torch.nn.BatchNorm2d
Figure 4: Probing different PyTorch layer’s for their awareness of parameter differentiability. (3(a)) PyTorch’s nn.Linear is agnostic to parameter differentiability. (3(b), 3(c), 3(d)) PyTorch’s nn.ConvNd, (4(a), 4(b), 4(c)) nn.ConvTransposeNd, and (4(d)) nn.BatchNorm2d in evaluation mode are not agnostic to parameter differentiability.
Training mode + torch.compile
Evaluation mode + torch.compile
Figure 6: Probing different PyTorch layer’s for their awareness of parameter differentiability with torch.compile.
Appendix C Diagrams showing the computation graph for elementary layers
In all the following figures, we show PyTorch’s behavior of saving tensors when they are not required. For all layers, the Input case is shown.
The colors mean the following (taken from the torchviz package Zagoruyko et al. (2019)):
  • blue
    A node representing the main operation being discussed
  • green
    A node representing any tensor
  • gray
    A node representing any other operations (i.e. View/ AccumulateGrad etc.)
Tensors which are saved by the AD engine are marked as [saved tensor], and they also have an undirected edge to the main operation node.
C.1 Conv2d
PyTorch Conv2d - Input Case
Differentiable: Weights $\times$, Input ✓
Figure 7: Computation graphs of a convolution layer for the Input case. Even though the input is differentiable, PyTorch saves it (as can be seen inside the ConvolutionBackward0 node - the input is of shape [7, 3, 12, 12]). MemSave on the other hand, does not save the input.
C.2 Linear
PyTorch Linear - Input Case
Differentiable: Weights $\times$, Input ✓
Figure 8: Computation graphs of ${\bm{y}}={\bm{W}}{\bm{x}}+{\bm{b}}$ for the Input case. This visualization is interesting because these graphs indicate that a linear layer does not save its input if the weight is marked non-differentiable, which is already the optimal behaviour.
C.3 BatchNorm2d
Training Mode
PyTorch BatchNorm - Input Case
Differentiable: Weights $\times$×Input ✓
Figure 9: Computation graphs of ${\bm{y}}={\bm{W}}\frac{{\bm{x}}-{\bm{\mu}}({\bm{x}})}{\sqrt{{\bm{\sigma}}^{2}({\bm{x}})+\epsilon}}+{\bm{b}}$ (i.e., BatchNorm in training mode) for the Input case.
Evaluation Mode
PyTorch BatchNorm (eval mode) - Input Case
Differentiable: Weights $\times$×Input ✓
Figure 10: Computation graphs of ${\bm{y}}={\bm{W}}\frac{{\bm{x}}-\hat{{\bm{\mu}}}}{\sqrt{\hat{{\bm{\sigma}}}^{2}+\epsilon}}+{\bm{b}}$ (i.e., BatchNorm in eval mode) for the Input case. PyTorch also saves the input, even though it is not required for calculating the input gradient. MemSave recognizes this and does not save the input.